OpenStudy (anonymous):

What is the 9th term of the geometric sequence 4, 12, 36, …?

OpenStudy (anonymous):

\[\text{First term}\text = 4 \times 3^0\] \[\text{Second term} = 4 \times 3^1\] \[\text{nth term} \text = 4 \times 3^{n-1}\]